40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 7 18:20:01 2011, kiwamu, Summary, Green Locking, plans 
    Reply  Tue Mar 8 15:33:09 2011, kiwamu, Summary, Green Locking, plan on Mar.8th 
Message ID: 4387     Entry time: Tue Mar 8 15:33:09 2011     In reply to: 4382
Author: kiwamu 
Type: Summary 
Category: Green Locking 
Subject: plan on Mar.8th 
Today's goal is to measure the contribution from the intensity noise to the beatnote.
 
Plans for today
  - check the ADC for the DCPD that Jenne installed yesterday
  - adjust RF power on the AOM
  - take spectrum of the differential noise and measure the coupling from the intensity noise
  - update the noise budget

Quote: from #4382
This week's goal is to investigate the source of the differential noise and to lower it.

 

ELOG V3.1.3-